One angle of a triangle measures 6 degrees more than a second angle. The third angle measures 18 less than twice the sum of the first two angles. Find the measure of each angle.
Answer by mananth(16946)

One angle of a triangle measures 6 degrees more than a second angle. The third angle measures 18 less than twice the sum of the first two angles. Find the measure of each angle.
...
let second angle be x
first angle = x+6
third angle = 2(x+x+6)-18= 4x-6
...
x+x+6+4x-6=180
6x=180
/6
x=30
...
30, 36, 114 degrees
...
m.ananth@hotmail.ca